Milano Pizza is a small neighborhood pizzeria that has a small area for in-store dining as well as offering take-out and free home delivery services. The pizzeria’s owner has determined that the shop has two major cost drivers—the number of pizzas sold and the number of deliveries made. The pizzeria’s cost formulas appear below: Fixed Cost per Month Cost per Pizza Cost per Delivery Pizza ingredients $ 4.30 Kitchen staff $ 6,110 Utilities $ 710 $ 0.30 Delivery person $ 3.10 Delivery vehicle $ 730 $ 1.30 Equipment depreciation $ 480 Rent $ 2,070 Miscellaneous $ 830 $ 0.15 In November, the pizzeria budgeted for 1,860 pizzas at an average selling price of $17 per pizza and for 240 deliveries. Data concerning the pizzeria’s actual results in November appear below: Actual Results Pizzas 1,960 Deliveries 220 Revenue $ 33,970 Pizza ingredients $ 9,010 Kitchen staff $ 6,050 Utilities $ 935 Delivery person $ 682 Delivery vehicle $ 1,006 Equipment depreciation $ 480 Rent $ 2,070 Miscellaneous $ 850 Required: 1. Complete the flexible budget performance report that shows both revenue and spending variances and activity variances for the pizzeria for November. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)
Answer #1

Answer-

MILANO PIZZA
Flexible Budget Performance Report
For the month ended November
Revenue & Spending Variance Activity Variance
Particulars Variances Remark Variances Remark
$ $
Revenue 650 Favorable 1700 Favorable
Less:-Expenses
Pizza ingredients -582 Unfavorable -430 Unfavorable
Kitchen staff 60 Favorable 0 None
Utilities 363 Favorable -30 Unfavorable
Delivery person 0 None 62 Favorable
Delivery vehicle 10 Favorable 26 Favorable
Equipment 0 None 0 None
Rent 0 None 0 None
Miscellaneous 274 Favorable -15 Unfavorable
Total Expenses 125 Favorable -387 Unfavorable
Income from operations 775 Favorable 1313 Favorable

Explanation-

MILANO PIZZA
Revenue & Spending Variances
For the month ended November
Particulars Flexible Budget Actual Results Variances Remark
$ $ $
Revenue 1960 pizzas*$17 per pizza 33320 33970 650 Favorable
Less:-Expenses
Pizza ingrediants 1960 pizzas*$4.30 per pizza 8428 9010 -582 Unfavorable
Kitchen staff 6110 6050 60 Favorable
Utilities $710+(1960 pizzas*$0.30 per pizza) 1298 935 363 Favorable
Delivery person 220 deliveries*$3.10 per delivery 682 682 0 None
Delivery vehicle $730+(220 deliveries*$1.30 per delivery) 1016 1006 10 Favorable
Equipment 480 480 0 None
Rent 2070 2070 0 None
Miscellaneous $830+(1960 pizzas*$0.15 per pizza) 1124 850 274 Favorable
Total expenses 21208 21083 125 Favorable
Income from operations 12112 12887 775 Favorable
MILANO PIZZA
Activity Variance
For the month ended November
Particulars Flexible Budget Planning Budget Variances Remark
$ $
Revenue 1960 pizzas*$17 per pizza =33320 1860 pizzas*$17 per pizza = 31620 1700 Favorable
Less:-Expenses
Pizza ingredients 1960 pizzas*$4.30 per pizza= 8428 1860 pizzas*$4.30 per pizza= 7998 -430 Unfavorable
Kitchen staff 6110 6110 0 None
Utilities $710+(1960 pizzas*$0.30 per pizza) =1298 $710+(1860 pizzas*$0.30 per pizza) =1268 -30 Unfavorable
Delivery person 220 deliveries*$3.10 per delivery=682 240 deliveries*$3.10 per delivery=744 62 Favorable
Delivery vehicle $730+(220 deliveries*$1.30 per delivery) =1016 $730+(240 deliveries*$1.30 per delivery) =1042 26 Favorable
Equipment 480 480 0 None
Rent 2070 2070 0 None
Miscellaneous $830+(1960 pizzas*$0.15 per pizza)=1124 $830+(1860 pizzas*$0.15 per pizza)=1109 -15 Unfavorable
Total Expenses 21208 20821 -387 Unfavorable
Income from operations 12112 10799 1313 Favorable